A small piston of a hydraulic press has an area of 20.0 cm2. If the applied force to the piston is

75.0 N, what must the area of the connected large piston be to exert a force of 6550 N?


1
Expert's answer
2021-01-19T07:08:21-0500

Given:

"A_1=20.0\:\rm cm^2"

"F_1=75.0\:\rm N"

"F_2=6550\:\rm N"

The Pascal's principle says

"\frac{F_1}{A_1}=\frac{F_2}{A_2}"

Hence

"A_2=A_1\frac{F_2}{F_1}=\rm 20.0\times\frac{6550}{75.0}=1750\: cm^2"

Answer: "\rm 1750\: cm^2"